You are given the overall reliability in a two-component parallel system is 0.979. Calculate the reliability of the second component if the reliability of the initial component is 0.97.

Group of answer choices

0.40

0.30

0.25

0.35
Answer #1

Given:   Reliability of two component parallel system is 0.979.

               Reliability of the initial component (say R1) is 0.97

Solution:

Reliability, R, is given by the formula:

R = 1 - F, where F is the probability of failure and F = F1* F2* .... *Fn

As per the given data:

R = 0.979 = 1 - F

→ F = 1 - 0.979

F = 0.021

&

R1 = 0.97 = 1 - F1

→ F1 = 1-0.97

F­1 = 0.03

Considering a two component parallel system,

F = F1 * F2

→ 0.021 = 0.03 * F2

F2 = 0.7

R2 = 1 - F2

→ R2 = 1 - 0.7

R2 = 0.30

Reliability of the second component is 0.30.